### v1

# Reciprocal Reward Influence Encourages Cooperation From Self-Interested Agents

Abstract
Emergent cooperation among self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents. Instead, naïve reinforcement learning algorithms typically converge to Pareto-dominated outcomes in even the simplest of social dilemmas. An emerging class of opponent-shaping methods have demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents. However, they rely on higher-order derivatives through the predicted learning step of other agents or learning meta-game dynamics, which in turn rely on stringent assumptions over opponent learning rules or exponential sample complexity, respectively. To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of an opponent’s actions on their returns. This approach effectively seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without attempting to directly shape policy updates. We show that Reciprocators can be used to promote cooperation in a variety of temporally extended social dilemmas during simultaneous learning.
1 Introduction
Many species exhibit cooperative behaviors of remarkable variety and complexity. Even among prosocial animals, however, humans are especially notable for their ability to cooperate with unrelated individuals and maintain that cooperation even in highly adversarial environments (Melis & Semmann, 2010). These qualities are often credited with the development of human technology, culture, and advanced cognition (Burkart et al., 2014). As artificially intelligent (AI) agents are integrated into human society, it is increasingly important to ensure that they share these same prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
However, state-of-the-art reinforcement learning (RL) methods are typically designed for the single agent setting and are ill-suited for the nonstationarities introduced by multiple agents learning simultaneously. Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce globally suboptimal outcomes. This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tensions between collective and individual return (Leibo et al., 2017). SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for fully-cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018). Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs, including mutual confession in the iterated Prisoners’ Dilemma (IPD) and an expected reward of 0 in Coins (Foerster et al., 2018).
In this work, we use an intrinsic reward to encourage agents called Reciprocators to reciprocate the influence that others exert on their returns, with the goal of inducing mutually beneficial behavior from other agents. We first define a notion of influence that quantifies the effect that one agent’s action has on another’s expected return. Then, given a pair of agents, a Reciprocator $rc$ and another agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we call the influence balance that agent $i$ owes $rc$. At any time step, the Reciprocator receives an intrinsic reciprocal reward proportional to the product of its influence balance against agent $i$ and the influence of its selected action on agent $i$’s return. This incentivizes the Reciprocator to take reciprocating actions whose influence matches the sign of its current influence balance. For example, if agent $i$ has a net positive influence balance with $rc$, then $rc$’s intrinsic reciprocal reward will be positive when taking actions that positively affect agent $i$’s return.
The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation by influencing others’ expected returns without modifying the reward structure of the environment, requiring higher-order derivatives, or learning meta-game dynamics. (2) We show that agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas without needing training schedules to control opponent selection. (3) Agents equipped with our intrinsic reward are able to induce cooperative behavior from purely self-interested naïve learners and several higher-order baselines.
2 Related Work
In order to improve convergence towards Pareto-optimal solutions among independently learning agents, previous work has modified agents’ reward structures to explicitly consider either group or per-capita return in order to promote cooperative behavior, e.g., via inequity aversion (Hughes et al., 2018), “empathic” harm reduction (Bussmann et al., 2019), or “altruistic” gradient adjustments (Li et al., 2024). While agents that abide by such restrictions may seek to reduce their harmful externalities, they have no way of regulating the behavior of other, less magnanimous agents, including purely self-interested “exploiters” in the worst case (Agapiou et al., 2023). Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong assumptions over the other agents’ altruistic tendencies.
Another avenue towards a more robust form of cooperation relies on the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010). Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of sequential social dilemmas (SSDs). However, these methods require extensions of the original action space that allow agents to directly modify other agents’ rewards. On the other hand, we instead propose to quantify the influence of each naturally available action $a\in A$ in a given state $s$ on another agent’s expected return $R$, and use this influence as the medium for reciprocation. As we will show, such reciprocal reward and punishment can be seen as a form of “opponent shaping” that seeks to manipulate the $Q$-value of given state-action pair and consequently the likelihood of that action in future policies.
Considering future policies points to a key issue with the canonical reinforcement learning (RL) framework, in that it optimizes only for the expected return within a single episode of the environment. However, taking actions that seek to optimize the long-term behavior of the other agents in the environment will not receive any immediate feedback within an episode, and must wait for one, or possibly many, learning steps. Opponent shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020, LIO), through opponents’ gradient updates (Foerster et al., 2018, LOLA), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022, MFOS). While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity to learn the dynamics of meta-games, respectively.
In particular, LOLA and subsequent works such as Proximal Learning with Opponent-Learning Awareness (Zhao et al., 2022, POLA) and Consistent Learning with Opponent-Learning Awareness (Willi et al., 2022, COLA) differentiate through the opponent’s learning update but have only been demonstrated to work with vanilla gradient descent and fixed learning rates. However, modeling the learning rule of each opponent becomes increasingly implausible in the face of modern RL algorithms, which uses adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016). Indeed, LOLA with opponent modeling (LOLA-OM) showed significantly worse results than agents with direct access to their opponent’s learning rule. Another class of opponent-shaping algorithms such as MFOS uses meta-learning approaches, which must repeatedly test new meta-policies in full training runs known as “meta-episodes” in order to learn how to exploit opponent learning dynamics. This requires the ability to freely perform rollouts in the environment against opponents whose policies can be reset to initialization at will, an unrealistic assumption in environments with partially adversarial motivations.
Our work is conceptually similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners. However, LOQA still differentiates through a model of the opponent’s policy and optimizes according to the joint advantage function $A(s_{t},a_{1},a_{2})$ computed with respect to the state-value function $V(s_{t})$. On the other hand, we use a counterfactual baseline that marginalizes out only the opponent’s action to perform agent-specific credit assignment and use an intrinsic reward instead of gradients over opponent policies to encourage agents to influence the opponent’s $Q$-values in the correct direction. Most importantly, LOQA focuses on the problem of learning a general end-policy that performs well against a variety of other agents at evaluation and has therefore only demonstrated the ability to shape other LOQA opponents in a controlled self-play scenario, while our method focuses on shaping the policies of diverse agents over multiple episodes of simultaneous learning.
In the context of these higher-order opponent-shaping approaches, we position our intrinsic reciprocal reward as a form of immediate, within-episode feedback that encourages otherwise naïve learners to implicitly consider the long-term, cross-episode effects of their actions on the policy changes of other agents, demonstrating many of the properties of higher-order opponent-shaping methods while using only first-order reinforcement learning algorithms and standard rollout-based training procedures.
3 Preliminaries
We construct a series of sequential social dilemmas which can each be described as a stochastic game $G$, defined by a tuple $G=\langle S,A,P,r,n,\gamma\rangle$. In $G$, $n$ agents, indexed by $i\in\{1,\ldots,n\}$, observe the state of the environment $s\in S$ and simultaneously choose actions $a^{i}\in A$ to form a joint action $\boldsymbol{a}\in\boldsymbol{A}\equiv A^{n}$. The environment then undergoes a change according to the state transition function $P(s^{\prime}|s,\boldsymbol{a}):S\times\boldsymbol{A}\times S\rightarrow[0,1]$. Each agent receives an individual reward $r^{i}=r^{i}(s,\boldsymbol{a})$, and future rewards are discounted at each time step by a discount factor $\gamma\in[0,1]$. We consider the fully observable setting where agents observe the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent. Each agent $i$ conditions a stochastic recurrent policy $\pi^{i}(a^{i}|\tau^{i})$ on its action-observation history, which is denoted as $\tau^{i}\in T\equiv(S\times\boldsymbol{A})^{*}$.
4 Reciprocal Reward Influence
4.1 1-Step Value Influence
Wang et al. (2020) proposed an intrinsic reward for exploration called Value of Interaction (VoI). VoI measures how much agent $i$’s actions affect agent $j$’s expected return by comparing the state-action value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ against a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$, where the $\boldsymbol{\pi}$ superscript indicates the dependence of this function on the parameters of all agents’ policies. This baseline marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return. This difference of $Q$-values can be decomposed into an immediate influence $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence computed over changes in state transition probabilities [Equation 2]. Although the original VoI was formulated as a regularizer using an expectation over trajectories $\tau$, we modify it to compute the one-step Value Influence (VI) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to compute the influence of individual actions
$VI_{i\mid j}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}\right)$ $=Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)-Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$ (1)
$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\boldsymbol{\pi}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}).$ (2)
This definition of influence relies on the notion of a counterfactual baseline to assign credit to a particular agent’s action while holding all other agents’ actions constant. This counterfactual baseline is so named because it estimates the counterfactual expected return if the agent’s true action is replaced with a “default” action, which is computed by marginalizing out the agent’s action to get the expected on-policy return. This baseline return is given by
$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
In practice, we approximate this by regressing towards the Q-value while masking out $a_{i}$ from the joint state-action input. The concept of marginalizing out an agent’s action as a baseline to quantify influence has also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $i$’s expected reward that can be achieved by agent $j$’s action by taking the minimum over $a^{j}\in A$ in Equation 1. Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus, which is a special case of $VI_{i|j}$ where $i=j$ (Equation 4).
$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)
4.2 Keeping Score with Influence Balances
The amount of influence that a Reciprocator is able to exert in a single time step is heavily environment-dependent, so that it may not always be possible to immediately reciprocate previous influences. To encourage reciprocation over extended timescales, we continuously accumulate a measure of net influence over sequences of actions. We draw inspiration from the “debit” mechanism $D_{i}$ used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT) to accumulate the advantage $A$ gained by an opponent $i$ compared to a cooperative counterfactual baseline, which they compute by recursively summing $A$ at every timestep.
Extending this idea to our framework, we sum agent $i$’s influence $VI^{\boldsymbol{\pi}}_{i|rc}$ on the Reciprocator’s expected return at each timestep rather than its own. However, using influence in only one direction as motivation for reciprocation can lead to continuous punishment or rewarding without a way to settle the score. To mitigate this, we also accumulate the net influence $VI^{\boldsymbol{\pi}}_{rc|i}$ in the opposite direction and subtract it from the influence balance at every timestep as a way to “pay off” the balance, limiting the degree to which reciprocation is rewarded. Formally, we define the influence balance $B_{rc|i}(t)$ maintained by a Reciprocator $rc$ with another agent $i$ at time $t$ as
$B_{rc|i}(t)=B_{rc|i}(t-1)+[VI^{\boldsymbol{\pi}}_{i|rc}(s_{t},\boldsymbol{a}_{t})-VI^{\boldsymbol{\pi}}_{rc|i}(s_{t},\boldsymbol{a}_{t})].$ (5)
The influence balance can be thought of as a score of net influence over time between agents, and can be used to motivate reciprocation in the correct direction, i.e., either positive reinforcement of net positive influence or positive punishment of net negative influence.
4.3 Intrinsic Reciprocal Reward
If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over the learned baseline expectation, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates. Similar logic applies for detrimental deviations, negative influence balance, and subsequent reciprocal punishment. We then define the intrinsic reciprocal reward $r^{R}_{rc|i}(t)$ received by $rc$ as
$r^{R}_{rc|i}(t)=B_{rc|i}(t-1)\cdot VI_{rc\mid i}^{\boldsymbol{\pi}}\left(s_{t},\boldsymbol{a}_{t}\right).$ (6)
Taking the product of existing influence balance and current action’s VI encourages the Reciprocator to take actions that reinforce agent $i$’s behavior in the correct direction by matching signs, and scales the reward by the magnitude of the outstanding influence balance and the reinforcing influence. The intrinsic reward is then added to the agent’s extrinsic reward to form the total reward used in training.
Algorithm 1 Training with Reciprocal Reward Influence vs. Agent $i$
  Initialize agent $rc$’s policy parameters $\theta_{\pi}$, VI target function parameters $\phi$, influence balance vector $B_{rc|i}=\boldsymbol{0}$, policy memory $\mathcal{M}=\emptyset$, and influence memory $\mathcal{H}=\emptyset$
  for each episode $e$ do
     Observe initial state $\boldsymbol{s_{0}}$
     for $t=1$ to $T$ do
        Choose action $a_{t}\sim\pi_{\theta}(s_{t})$
        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},\boldsymbol{s}_{t+1}$
        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(\boldsymbol{s_{t}},\boldsymbol{a_{t}},\boldsymbol{r_{t}},\boldsymbol{s_{t+1}})$ in $\mathcal{H}$
     end for
     Update influence target function parameters $\phi$ with $\mathcal{H}$ every $k$ episodes
     Compute reciprocal rewards $r^{R}_{rc|i}(1),\ldots,r^{R}_{rc|i}(T)$ w.r.t. agent $i$ and sum with rewards in $\mathcal{M}$
     Compute advantage estimates $\hat{A}_{1},\ldots,\hat{A}_{T}$
     for $K$ epochs do
        Optimize the surrogate PPO-clip objective w.r.t. $\theta_{\pi}$
     end for
     Reset $\mathcal{M}=\emptyset$
  end for
4.4 Policy Optimization with Reciprocal Rewards
We use experience replay to periodically update target networks and iteratively update our counterfactual baseline estimates towards these target values (Mnih et al., 2015). This provides two key benefits: first, periodically updating these policy-dependent functions stabilizes training and allows us to approximately ignore their gradients, and therefore the gradient of the intrinsic reciprocal reward, with respect to the agents’ policy parameters (Wang et al., 2020). With this assumption, we are able to use to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
Second, drawing samples from multiple previous episodes to the train the counterfactual baseline target functions makes the Reciprocator less susceptible to exploitation. If the counterfactual estimators were updated concurrently with the policy using only the most recent on-policy data, then the Reciprocator’s baseline would be immediately adjusted to its opponent’s new policy after each episode. Because assessment of influence hinges on these counterfactual baselines, updating them too frequently would allow adversaries to easily manipulate these estimates of on-policy returns.
5 Experiments
We conduct experiments using two commonly-used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents. For IPD, we consider memory-1 iterated games as in Foerster et al. (2018) and Lu et al. (2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories. We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al. (2018), and more standard batched policy rollouts for a fixed episode length. We append “analytic” and “rollout” to the game names to denote the evaluation method used. For Coins, we augment the observation with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022).
5.1 Sequential Social Dilemmas
Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally-extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a). The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
Coins: Coins is an $n$-player game introduced by Lerer & Peysakhovich (2018) as a temporally-extended version of the IPD. In this game, $n$ players move around in a fixed-size grid and collect colored coins. Agents receive a reward of 1 for collecting a coin, and are punished if another agent collects their coin such that, if agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0. We show example rewards for $n=2$ in Figure 1(b).
C D
C -1, -1 -3, 0
D 0, -3 -2, -2
(a) Rewards for IPD.
(b) Rewards for Coins.
Figure 1: (a) The first number in each cell denotes the reward received by the agent taking the row action, and the second the reward received by the agent taking the column action, where C: cooperate (stay silent) and D: defect (confess). (b) Two agents (red and blue) are tasked with collecting randomly spawning coins. If an agent collects its own coin, it receives a reward of +1 (left). If an agent collects another’s coin, then it receives a reward of +1 but the other agent receives a punishment of -2.
5.2 Baselines
Although we have discussed an extensive body of work related to cooperation in SSDs, we select these baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning.
Naïve Learner (NL): As previously defined, NLs optimize their expected return with respect only to their own policy parameters $\theta$. In this work, we implement NLs using standard policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
Table 1: IPD-Analytic round robin results
RC NL LOLA M-MAML MFOS
RC -1.06 -1.03 -1.05 -1.05 -1.06
NL -1.06 -1.98 -1.52 -1.28 -1.88
LOLA -1.08 -1.30 -1.09 -1.04 -1.02
M-MAML -1.13 -1.25 -1.15 -1.17 -1.56
MFOS -0.98 -0.65 -1.02 -0.81 -1.01
Each entry is the average reward per episode achieved by the row agent against the column agent. We bold cooperative outcomes, which we define as both agents’ outcomes being better than -1.25, with a difference of no more than 0.10 between them. Standard error of the mean is less than 0.01 for all experiments. M-MAML results are averaged across 10 initial policies.
Figure 2: Representative run of a Reciprocator vs. an NL in IPD-Rollout. Reciprocal reward per step (left axis) and percentage of cooperation (staying silent) over the course of an episode (right axis).
Learning with Opponent-Learning Awareness (LOLA): LOLA uses either whitebox access to an opponent’s gradients and Hessians or an explicit model of their policy parameters, assuming they are NLs, and differentiates through their learning step using the update rule
$\theta_{t+1}^{i}$ $=\theta_{t}^{i}+\alpha^{i}\nabla_{\theta_{t}^{i}}J^{i}\left(\theta_{t}^{i},\theta_{t}^{-i}+\Delta\theta_{t}^{-i}\right)$
$\Delta\theta_{t}^{-i}$ $=\alpha^{-i}\nabla_{\theta_{t}^{i}}J^{-i}\left(\theta_{t}^{i},\theta_{t}^{-i}\right).$
Multiagent Model-Agnostic Meta-Learning (M-MAML): M-MAML (Lu et al., 2022) learns initial parameters and then meta-learns over both its own and its opponent’s policy updates, conceptually similar to Meta-Multiagent Policy Gradient (Kim et al., 2021, Meta-MAPG) but modified to differentiate directly through the analytic form of the return in matrix games.
Model-Free Opponent Shaping (MFOS): As briefly discussed in the introduction, MFOS meta-learns over multiple episodes of policy updates in order to accomplish long-horizon opponent shaping. In Lu et al. (2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is elementwise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
5.3 Implementation Details
In IPD-Analytic, we differentiate directly through the analytic solution to the matrix game. For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization and a clipped surrogate objective (Schulman et al., 2017, PPO-Clip). Target networks to estimate the $Q$-values in Equation 1 are updated every $k$ episodes using uniformly sampled experience from a replay buffer (Lin, 1992). Additional hyperparameter values and network architecture details can be found in Appendix A.1. Shaded regions in plots depict one standard error of the mean.
6 Results
Here we summarize the results of our experiments, which include round-robin evaluations in IPD and demonstrations of Reciprocators’ ability to reach cooperation and shape NLs in Coins. Following work by Yu et al. (2022) showing that Meta-MAPG and LOLA-DICE agents fail to achieve significant results in Coins, even with a simplified shared reward, we do not use them as baselines in Coins.
6.1 IPD
IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines despite being only a first-order method using vanilla gradient descent [Table 1]. Although the Reciprocator is extorted by MFOS to small extent (-1.06 vs. -0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs and roll out new meta-policies against opponents that are repeatedly reset back to randomly initialized policies. Additionally, as explained in Section 4, stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts.
IPD-Rollout: In IPD-Rollout, we show that Reciprocator is able to consistently shape an NL in IPD over the course of a single learning trajectory with limited rollout samples. We observe interesting learning dynamics wherein the Reciprocator’s policy repeatedly oscillates between lower and higher levels of cooperation. We show that these oscillations are driven by the opposing intrinsic and extrinsic rewards, and observe that the derivative of the rate of cooperation corresponds to the reciprocal reward [Figure 2]. Mechanistically, we see that the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward. However, as the NL’s policy becomes more deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases. As the extrinsic reward begins to dominate, the Reciprocator essentially reverts into an NL, leading to exploitative defection. However, this exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term to become negative, leading to a negative reciprocal reward. This, combined with an increase in the NL’s frequency of defection (leading to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$), produces a reversal back to cooperative behavior that is then reinforced by positive reciprocal rewards.
Figure 3: Shaping an NL in Coins. Percentage of own coins collected by NL during training when facing each opponent (left) and coin counts across 32 steps for Reciprocator vs. NL (right).
6.2 Coins
In this temporally-extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates equal to MFOS. This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to this behavior (32,000 steps for Reciprocator vs. 512,000 steps for MFOS) [Figure 3]. We show that this change is driven by changes in coin preference rather than in total collection, with both agents collecting at near-optimal pace. This is in contrast to MFOS [Appendix B.2], which shapes the NL into collecting fewer coins.
When two Reciprocators are pitted against each other, we see that they are quickly able to learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play. This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieved an average reward of only $\sim 2$ per 32 steps according to Foerster et al. (2018). One concern is that the intrinsic reciprocal reward overpowers the extrinsic reward, effectively changing the motivational structure of the game into one that is fully-cooperative. To address this, we additionally recorded the cumulative reciprocal reward received in each episode over the course of training [Appendix B.1] to show that the reciprocal reward quickly decreases to a negliglible fraction of the total reward.
Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind). Total reward per episode (left), percentage of own coins collected (right). Reciprocators and Naive Learner results are plotted on a scale of single episodes (bottom x-axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode is equivalent to 16 episodes (top x-axis).
Together, these results demonstrate that Reciprocators are able to robustly shape the behavior of other agents towards prosocial equilibria during simultaneous learning, achieving state-of-the-art results with fewer assumptions and limitations than existing methods. Apart from opponent-shaping properties, we also show that the intrinsic reciprocal reward discourages Reciprocators from exploiting others, showing promise for the development of a more cooperative multi-agent learning framework.
7 Limitations and Future Directions
Our current implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes. Using handpicked weights to balance intrinsic and extrinsic rewards opens the door for a suboptimal tradeoff between long-term opponent shaping and short-term return maximization. This is partially addressed by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2]. Future work will focus on methods to evaluate reciprocation efficacy across multiple episodes and dynamically tune the balance between reciprocal and extrinsic rewards.
8 Conclusion and Broader Impacts
Emerging interest in cooperative AI has been led by approaches that endow agents with higher-order shaping capabilities. Although these agents exhibit cooperative behaviors when pitted against equivalent opponents, they readily manipulate and exploit agents with simpler learning rules or lower computational capabilities. This is a fundamentally undesirable outcome in partially-adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box performance.
We presented Reciprocators, naïve learning agents equipped with an intrinsic reward for multi-agent settings that seeks to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes. We showed that Reciprocators are able to shape the behavior of other agents in the environment over the course of training, and can produce prosocial behaviors in a variety of sequential social dilemmas. To the best of our knowledge, Reciprocators comprise the first class of learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or training schedules such as self-play or tracing procedures which control opponent selection and rely on the symmetric player roles. We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning algorithms.
Appendix A Appendix / supplemental material
All experiments were run on an Nvidia 3070 GPU with 8 GB of VRAM. For IPD-Analytic and Coins experiments involving MFOS agents, we adapted the original code from Lu et al. (2022).
A.1 Hyperparameter Details
Each experiment was run 8 times, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins.
Hyperparameter IPD-Rollout Coins
Number of Linear Layers 2 2
Size of Linear Layers 2 16
Number of GRUs - 1
Size of GRUs - 16
Episode Length 32 32
Adam Learning Rate 0.005 0.005
PPO Epochs Per Episode $K$ 10 40
PPO-Clip $\epsilon$ 0.1 0.15
Discount Factor $\lambda$ 0.96 0.99
Entropy Coefficient 0.02 0.01
Table 2: General PPO parameters.
With the exception of a linear layer size of 32 for Coins, network parameters to estimate the various target functions to compute the VI are the same as those listed for the PPO network in Table 2. Target estimates for IPD policies were computed by averaging across observed choices for each of the five possible states (start, CC, CD, DC, DD). Replay buffer sizes are in units of episodes, where each episode consists of a batch of 32 steps. Then, for example, a replay buffer size of 4 for Coins corresponds to 32 steps $\times$ 4 episodes $\times$ 2048 batch size = 262,144 steps of experience stored in the replay buffer.
Hyperparameter IPD-Analytic IPD-Rollout Coins
Replay Buffer Size (in episodes) 5 1 4
Target Function Update Period 10 3 1
Target Function Epochs Per Episode - - 20
Adam Learning Rate - - 0.01
Reciprocal Reward Weight 5.0 5.0 1.0
Table 3: Reciprocator-specific parameters.
Again, network parameters for the outer meta-policy and inner actor and critic networks are the same as the general PPO parameters above.
Hyperparameter IPD-Analytic IPD-Rollout Coins
Outer Episode Length 16 16 16
Inner Episode Length - 32 32
Table 4: MFOS-specific parameters.
Appendix B Additional Experiments
B.1 Reciprocal Reward in Coins
An example run of two Reciprocators in Coins showing the relative magnitude of intrinsic reciprocal reward and extrinsic environmental rewards over the course of training.
Figure 5: Mean extrinsic reward (left), mean intrinsic reward (right) for two Reciprocators in Coins.
B.2 Coin Counts vs. NL-PPO
We show the total coin counts over the course of training runs vs. an NL-PPO in order to show that changes in P(Own Coins) are driven by changes in coin color preference rather than changes in the total number of coins collected.
Figure 6: Percentage of own coins collected by NL-PPO (left) and total number of coins collected by NL-PPO (right) vs. each baseline in Coins.

### v2

@@
 Abstract
-Emergent cooperation among self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents.
+Cooperation between self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents.
 Instead, naïve reinforcement learning algorithms typically converge to Pareto-dominated outcomes in even the simplest of social dilemmas.
-An emerging class of opponent-shaping methods have demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents.
-However, they rely on higher-order derivatives through the predicted learning step of other agents or learning meta-game dynamics, which in turn rely on stringent assumptions over opponent learning rules or exponential sample complexity, respectively.
-To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of an opponent’s actions on their returns.
-This approach effectively seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without attempting to directly shape policy updates.
-We show that Reciprocators can be used to promote cooperation in a variety of temporally extended social dilemmas during simultaneous learning.
+An emerging literature on opponent shaping has demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents.
+However, such methods differentiate through the learning step of other agents or optimize for meta-game dynamics, which rely on privileged access to opponents’ learning algorithms or exponential sample complexity, respectively.
+To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of opponents’ actions on their returns.
+This approach seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without directly differentiating through a model of their policy.
+We show that Reciprocators can be used to promote cooperation in temporally extended social dilemmas during simultaneous learning.
+Our code is available at https://github.com/johnlyzhou/reciprocator/.
 1 Introduction
@@
 These qualities are often credited with the development of human technology, culture, and advanced cognition (Burkart et al., 2014).
-As artificially intelligent (AI) agents are integrated into human society, it is increasingly important to ensure that they share these same prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
+As artificially intelligent (AI) agents become more commonplace in human society, it is increasingly important to ensure that they share similar prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
 However, state-of-the-art reinforcement learning (RL) methods are typically designed for the single agent setting and are ill-suited for the nonstationarities introduced by multiple agents learning simultaneously.
-Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce globally suboptimal outcomes.
-This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tensions between collective and individual return (Leibo et al., 2017).
-SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for fully-cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018).
-Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs, including mutual confession in the iterated Prisoners’ Dilemma (IPD) and an expected reward of 0 in Coins (Foerster et al., 2018).
-In this work, we use an intrinsic reward to encourage agents called Reciprocators to reciprocate the influence that others exert on their returns, with the goal of inducing mutually beneficial behavior from other agents.
-We first define a notion of influence that quantifies the effect that one agent’s action has on another’s expected return.
-Then, given a pair of agents, a Reciprocator $rc$ and another agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we call the influence balance that agent $i$ owes $rc$.
-At any time step, the Reciprocator receives an intrinsic reciprocal reward proportional to the product of its influence balance against agent $i$ and the influence of its selected action on agent $i$’s return.
-This incentivizes the Reciprocator to take reciprocating actions whose influence matches the sign of its current influence balance.
-For example, if agent $i$ has a net positive influence balance with $rc$, then $rc$’s intrinsic reciprocal reward will be positive when taking actions that positively affect agent $i$’s return.
-The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation by influencing others’ expected returns without modifying the reward structure of the environment, requiring higher-order derivatives, or learning meta-game dynamics.
-(2) We show that agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas without needing training schedules to control opponent selection.
-(3) Agents equipped with our intrinsic reward are able to induce cooperative behavior from purely self-interested naïve learners and several higher-order baselines.
+Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce collectively suboptimal outcomes.
+This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tradeoffs between collective and individual return (Leibo et al., 2017).
+SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018).
+Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs (Foerster et al., 2018).
+In this work, we propose an intrinsic reward that encourages agents called Reciprocators to learn a tit-for-tat-like strategy which reciprocates the influence that others exert on their returns.
+We first define value influence, a notion of influence that quantifies the effect that one agent’s action has on another’s expected return.
+Given a pair of agents, a Reciprocator $rc$ and another learning agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we refer to as the influence balance “owed” to $rc$ by agent $i$.
+At each time step, the Reciprocator receives an additional intrinsic reciprocal reward proportional to the product of its current influence balance with agent $i$ and the value influence of its action on agent $i$.
+This encourages the Reciprocator to take actions whose influence matches the sign of the influence balance: for example, if agent $i$ has produced a net positive influence on $rc$, then $rc$ will be rewarded for actions that have a reciprocally positive influence on agent $i$’s expected return.
+Mechanistically, our method seeks to manipulate the opponent’s $Q$-values by altering their expected return in particular directions depending on their actions.
+A reward-maximizing agent $i$ should then be incentivized to take mutually beneficial actions and avoid harmful externalities.
+The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation from other, simultaneously learning agents without modifying the structure of the environment, computing higher-order derivatives, or learning meta-game dynamics.
+(2) Agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas and are able to shape purely self-interested naïve learners into mutually beneficial behavior.
+(3) Reciprocators demonstrate resistance to exploitation by higher-order baselines, despite using only first-order reinforcement learning algorithms.
 2 Related Work
@@
 While agents that abide by such restrictions may seek to reduce their harmful externalities, they have no way of regulating the behavior of other, less magnanimous agents, including purely self-interested “exploiters” in the worst case (Agapiou et al., 2023).
-Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong assumptions over the other agents’ altruistic tendencies.
-Another avenue towards a more robust form of cooperation relies on the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010).
-Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of sequential social dilemmas (SSDs).
+Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong guarantees over the other agents’ altruistic tendencies.
+2.1 Cooperation through Influence
+A more robust form of cooperation can be achieved by actively exerting influence over other agents in the environment rather than unilaterally adopting a prosocial policy, especially in SSDs where extrinsic rewards incentivize defection.
+Jaques et al.
+(2019) provides an intrinsic reward for “social influence” on other agents, which is computed as the Kullback-Leibler (KL) divergence between an opponent’s policy distribution conditioned on an influence agent’s action $a^{i}$ and a counterfactual distribution which marginalizes out that influence agent.
+However, as a metric-agnostic quantity, KL divergence is unable to capture key details of how the action distribution changes (Park et al., 2024).
+In particular, this definition of influence cannot selectively modify opponents’ behavior with respect to extrinsic task returns or state transition dynamics, and therefore has limited utility as a mechanism to incentivize cooperation.
+Other works seek to directly influence opponent returns via the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010).
+Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of SSDs.
 However, these methods require extensions of the original action space that allow agents to directly modify other agents’ rewards.
 On the other hand, we instead propose to quantify the influence of each naturally available action $a\in A$ in a given state $s$ on another agent’s expected return $R$, and use this influence as the medium for reciprocation.
-As we will show, such reciprocal reward and punishment can be seen as a form of “opponent shaping” that seeks to manipulate the $Q$-value of given state-action pair and consequently the likelihood of that action in future policies.
+As we later show, selectively rewarding or punishing an opponent’s actions can be seen as a form of opponent shaping which seeks to manipulate the $Q$-value of given state-action pair in order to influence the likelihood of that action in future policies.
+2.2 Opponent Shaping
 Considering future policies points to a key issue with the canonical reinforcement learning (RL) framework, in that it optimizes only for the expected return within a single episode of the environment.
 However, taking actions that seek to optimize the long-term behavior of the other agents in the environment will not receive any immediate feedback within an episode, and must wait for one, or possibly many, learning steps.
-Opponent shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020, LIO), through opponents’ gradient updates (Foerster et al., 2018, LOLA), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022, MFOS).
-While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity to learn the dynamics of meta-games, respectively.
-In particular, LOLA and subsequent works such as Proximal Learning with Opponent-Learning Awareness (Zhao et al., 2022, POLA) and Consistent Learning with Opponent-Learning Awareness (Willi et al., 2022, COLA) differentiate through the opponent’s learning update but have only been demonstrated to work with vanilla gradient descent and fixed learning rates.
-However, modeling the learning rule of each opponent becomes increasingly implausible in the face of modern RL algorithms, which uses adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016).
-Indeed, LOLA with opponent modeling (LOLA-OM) showed significantly worse results than agents with direct access to their opponent’s learning rule.
-Another class of opponent-shaping algorithms such as MFOS uses meta-learning approaches, which must repeatedly test new meta-policies in full training runs known as “meta-episodes” in order to learn how to exploit opponent learning dynamics.
-This requires the ability to freely perform rollouts in the environment against opponents whose policies can be reset to initialization at will, an unrealistic assumption in environments with partially adversarial motivations.
-Our work is conceptually similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners.
+Opponent-shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020), through opponents’ gradient updates (Foerster et al., 2018; Zhao et al., 2022; Willi et al., 2022), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022).
+While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity (Fung et al., 2023) to learn the dynamics of meta-games, respectively.
+In particular, the Learning with Opponent-Learning Awareness (Foerster et al., 2018, LOLA) class of approaches differentiate through the opponent’s learning update but have only been demonstrated to work with full access to either the opponent’s policy gradients and Hessians or their policy parameters.
+When using gradient approximations and modeling the opponent’s policy, LOLA with opponent modeling (LOLA-OM) showed significantly worse results, even against vanilla policy gradients.
+Modeling efforts become even more implausible when faced with modern RL techniques, which employ adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016) to improve learning.
+Model-Free Opponent Shaping (Lu et al., 2022, MFOS) was introduced as a more general meta-learning approach that requires neither privileged access to nor inconsistent assumptions on opponent learning rules.
+However, MFOS’s meta-policies are trained across multiple training runs repeated in sequence, each of which are treated as a single “meta-episode,” in order to learn how to exploit opponent learning dynamics.
+This requires the ability to freely perform rollouts in the environment against opponents whose policies can be repeatedly reset to initialization, an unrealistic assumption in environments with partially adversarial motivations.
+Our work is conceptually most similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners.
 However, LOQA still differentiates through a model of the opponent’s policy and optimizes according to the joint advantage function $A(s_{t},a_{1},a_{2})$ computed with respect to the state-value function $V(s_{t})$.
@@
 Each agent receives an individual reward $r^{i}=r^{i}(s,\boldsymbol{a})$, and future rewards are discounted at each time step by a discount factor $\gamma\in[0,1]$.
-We consider the fully observable setting where agents observe the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent.
+We consider the fully observable setting where agents have access to the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent.
 Each agent $i$ conditions a stochastic recurrent policy $\pi^{i}(a^{i}|\tau^{i})$ on its action-observation history, which is denoted as $\tau^{i}\in T\equiv(S\times\boldsymbol{A})^{*}$.
@@
 4.1 1-Step Value Influence
-Wang et al.
-(2020) proposed an intrinsic reward for exploration called Value of Interaction (VoI).
-VoI measures how much agent $i$’s actions affect agent $j$’s expected return by comparing the state-action value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ against a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$, where the $\boldsymbol{\pi}$ superscript indicates the dependence of this function on the parameters of all agents’ policies.
-This baseline marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return.
-This difference of $Q$-values can be decomposed into an immediate influence $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence computed over changes in state transition probabilities [Equation 2].
-Although the original VoI was formulated as a regularizer using an expectation over trajectories $\tau$, we modify it to compute the one-step Value Influence (VI) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to compute the influence of individual actions
+In order to compute a general measure of value-based influence, we modify an intrinsic exploration reward proposed by Wang et al.
+(2020) called Value of Interaction (VoI).
+VoI measures how much agent $i$’s actions affect agent $j$’s expected return by computing the difference between the $Q$-value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ and a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$ which marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return [Equation 1].
+Here, the bolded $\boldsymbol{\pi}$ superscript indicates the dependence of these $Q$-functions on the parameters of all agents’ policies.
+The VoI can be decomposed into an immediate influence term $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence term computed over changes in state transition probabilities [Equation 2].
+Although the original VoI was formulated as an expectation over trajectories $\tau$ and used as a regularizer, we modify it to compute the one-step Value Influence ($VI$) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to quantify the influence of individual actions
 $VI_{i\mid j}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}\right)$ $=Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)-Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$ (1)
-$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\boldsymbol{\pi}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}).$ (2)
+$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\pi^{i}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}),$ (2)
+where $r(s,\mathbf{a}^{-i})=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[r(s,(\mathbf{a}^{-i},a^{i}))\right]$ and $p(s^{\prime}|s,\mathbf{a}^{-i})=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[p^{\pi^{i}}(s^{\prime}|s,(\mathbf{a}^{-i},a^{i}))\right]$.
 This definition of influence relies on the notion of a counterfactual baseline to assign credit to a particular agent’s action while holding all other agents’ actions constant.
@@
 This baseline return is given by
-$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
+$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\mathbb{E}_{a^{i}\sim\pi^{i}}\left[Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i}))\right]=\sum_{a_{i}}\pi^{i}(a_{i}|\tau^{i})Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
 In practice, we approximate this by regressing towards the $Q$-value while masking out $a_{i}$ from the joint state-action input.
-The concept of marginalizing out an agent’s action as a baseline to quantify influence has also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $i$’s expected reward that can be achieved by agent $j$’s action by taking the minimum over $a^{j}\in A$ in Equation 1.
-Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus, which is a special case of $VI_{i|j}$ where $i=j$ (Equation 4).
-$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)
+Marginalizing out an agent’s action to quantify influence is a common paradigm in MARL, having also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $j$’s expected reward that can be achieved by agent $i$ as the minimum $VI_{i|j}^{\pi}(s,\boldsymbol{a})$ over all $a^{i}\in A$.
+Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus.
+Note that the COMA advantage function [Equation 4] can be expressed a special case of $VI_{i\mid j}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}\right)$ where $i=j$.
+$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau^{i})Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)
 4.2 Keeping Score with Influence Balances
@@
 To encourage reciprocation over extended timescales, we continuously accumulate a measure of net influence over sequences of actions.
-We draw inspiration from the “debit” mechanism $D_{i}$ used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT) to accumulate the advantage $A$ gained by an opponent $i$ compared to a cooperative counterfactual baseline, which they compute by recursively summing $A$ at every timestep.
+We draw inspiration from the “debit” tally used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT), which tracks the cumulative advantage gained by an opponent compared to a fully cooperative baseline policy known a priori.
 Extending this idea to our framework, we sum agent $i$’s influence $VI^{\boldsymbol{\pi}}_{i|rc}$ on the Reciprocator’s expected return at each timestep rather than its own.
@@
 4.3 Intrinsic Reciprocal Reward
-If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over the learned baseline expectation, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates.
+If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over a baseline estimate, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates.
 Similar logic applies for detrimental deviations, negative influence balance, and subsequent reciprocal punishment.
@@
   for each episode $e$ do
-     Observe initial state $\boldsymbol{s_{0}}$
+     Observe initial state $s_{0}$
      for $t=1$ to $T$ do
         Choose action $a_{t}\sim\pi_{\theta}(s_{t})$
-        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},\boldsymbol{s}_{t+1}$
-        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(\boldsymbol{s_{t}},\boldsymbol{a_{t}},\boldsymbol{r_{t}},\boldsymbol{s_{t+1}})$ in $\mathcal{H}$
+        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},s_{t+1}$
+        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(s_{t},\boldsymbol{a_{t}},\boldsymbol{r_{t}},s_{t+1})$ in $\mathcal{H}$
      end for
@@
 This provides two key benefits: first, periodically updating these policy-dependent functions stabilizes training and allows us to approximately ignore their gradients, and therefore the gradient of the intrinsic reciprocal reward, with respect to the agents’ policy parameters (Wang et al., 2020).
-With this assumption, we are able to use to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
+With this assumption, we are able to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
 Second, drawing samples from multiple previous episodes to the train the counterfactual baseline target functions makes the Reciprocator less susceptible to exploitation.
@@
 Because assessment of influence hinges on these counterfactual baselines, updating them too frequently would allow adversaries to easily manipulate these estimates of on-policy returns.
-5 Experiments
-We conduct experiments using two commonly-used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents.
-For IPD, we consider memory-1 iterated games as in Foerster et al.
-(2018) and Lu et al.
-(2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories.
-We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al.
-(2018), and more standard batched policy rollouts for a fixed episode length.
-We append “analytic” and “rollout” to the game names to denote the evaluation method used.
-For Coins, we augment the observation with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022).
-5.1 Sequential Social Dilemmas
-Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally-extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a).
-The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
-Coins: Coins is an $n$-player game introduced by Lerer & Peysakhovich (2018) as a temporally-extended version of the IPD.
-In this game, $n$ players move around in a fixed-size grid and collect colored coins.
-Agents receive a reward of 1 for collecting a coin, and are punished if another agent collects their coin such that, if agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0.
-We show example rewards for $n=2$ in Figure 1(b).
 C D
@@
 If an agent collects another’s coin, then it receives a reward of +1 but the other agent receives a punishment of -2.
+5 Experiments
+We conduct experiments using two commonly used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents.
+For IPD, we consider memory-1 iterated games as in Foerster et al.
+(2018) and Lu et al.
+(2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories.
+We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al.
+(2018), and more standard batched policy rollouts for a fixed episode length.
+We append “analytic” and “rollout” to the game names to denote the evaluation method used.
+For Coins, we augment the observation given to the critics and value influence estimators with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022), but do not provide them as input to the policy networks.
+5.1 Sequential Social Dilemmas
+Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a).
+The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
+Coins: Coins is a temporally extended variant of the IPD introduced by Lerer & Peysakhovich (2018).
+In this game, two coins spawn randomly in a fixed-size grid, with each coin corresponding to one of the two players (designated by color matching).
+Each player moves around the grid and receives a reward of 1 for collecting any coin, and a punishment of $-1$ if the other agent collects their coin.
+If agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0.
+We show example rewards for various scenarios in Figure 1(b).
 5.2 Baselines
-Although we have discussed an extensive body of work related to cooperation in SSDs, we select these baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning.
+We select the following baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning and without modifications to the environment.
+In Coins, we exclude Meta-MAPG and LOLA-DICE as baselines following work by Yu et al.
+(2022) showing that neither method is able to achieve significant results, even with a simplified shared reward.
 Naïve Learner (NL): As previously defined, NLs optimize their expected return with respect only to their own policy parameters $\theta$.
-In this work, we implement NLs using standard policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
+In this work, we implement NLs using policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
 Table 1: IPD-Analytic round robin results
@@
 Each entry is the average reward per episode achieved by the row agent against the column agent.
-We bold cooperative outcomes, which we define as both agents’ outcomes being better than -1.25, with a difference of no more than 0.10 between them.
-Standard error of the mean is less than 0.01 for all experiments.
-M-MAML results are averaged across 10 initial policies.
-Figure 2: Representative run of a Reciprocator vs.
-an NL in IPD-Rollout.
-Reciprocal reward per step (left axis) and percentage of cooperation (staying silent) over the course of an episode (right axis).
+Standard error of the mean (SE) is less than 0.01 for all experiments and M-MAML results are averaged across the 10 initial policies provided by Lu et al.
+(2022).
+Figure 2: Representative run of a Reciprocator vs. an NL in IPD-Rollout. Average reciprocal reward per step (left axis) and probability of cooperation (right axis) over the course of an episode.
 Learning with Opponent-Learning Awareness (LOLA): LOLA uses either whitebox access to an opponent’s gradients and Hessians or an explicit model of their policy parameters, assuming they are NLs, and differentiates through their learning step using the update rule
@@
 In Lu et al.
-(2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is elementwise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
+(2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is element-wise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
 5.3 Implementation Details
 In IPD-Analytic, we differentiate directly through the analytic solution to the matrix game.
-For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization and a clipped surrogate objective (Schulman et al., 2017, PPO-Clip).
+For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization using a clipped surrogate objective (Schulman et al., 2017, PPO-Clip).
+Results including naïve learners trained with this algorithm are denoted by NL-PPO.
 Target networks to estimate the $Q$-values in Equation 1 are updated every $k$ episodes using uniformly sampled experience from a replay buffer (Lin, 1992).
-Additional hyperparameter values and network architecture details can be found in Appendix A.1.
-Shaded regions in plots depict one standard error of the mean.
+Additional hyperparameter values and network architecture details can be found in Appendix A.
+Figure 3: Shaping an NL in Coins.
+Proportion of own coins collected by NL during training when facing each opponent (left) and coin counts by type for Reciprocator vs.
+NL (right).
+Reciprocator and NL-PPO results are plotted on a scale of single episodes (bottom axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode contains 16 episodes (top axis).
 6 Results
-Here we summarize the results of our experiments, which include round-robin evaluations in IPD and demonstrations of Reciprocators’ ability to reach cooperation and shape NLs in Coins.
-Following work by Yu et al.
-(2022) showing that Meta-MAPG and LOLA-DICE agents fail to achieve significant results in Coins, even with a simplified shared reward, we do not use them as baselines in Coins.
+We evaluate the Reciprocators’ performance against a variety of baselines in a round-robin tournament for IPD-Analytic.
+In Coins, we assess opponent-shaping performance against NLs and against another agent of the same type, which we refer to as symmetric Coins.
+Shaded regions indicate standard error of the mean (SE) over eight random seeds.
+We do not perform full round-robin experiments in non-analytic environments for the following reasons: the Reciprocators’ complex and stochastic learning rule renders LOLA’s assumptions invalid and the extended time to convergence due to reciprocal behavior makes collecting meta-training episodes for MFOS computationally prohibitive.
 6.1 IPD
-IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines despite being only a first-order method using vanilla gradient descent [Table 1].
+IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines, resisting exploitation by higher-order methods such as LOLA and MFOS despite being only a first-order method using vanilla gradient descent [Table 1].
 Although the Reciprocator is extorted by MFOS to small extent (-1.06 vs.
--0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs and roll out new meta-policies against opponents that are repeatedly reset back to randomly initialized policies.
-Additionally, as explained in Section 4, stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts.
-IPD-Rollout: In IPD-Rollout, we show that Reciprocator is able to consistently shape an NL in IPD over the course of a single learning trajectory with limited rollout samples.
-We observe interesting learning dynamics wherein the Reciprocator’s policy repeatedly oscillates between lower and higher levels of cooperation.
-We show that these oscillations are driven by the opposing intrinsic and extrinsic rewards, and observe that the derivative of the rate of cooperation corresponds to the reciprocal reward [Figure 2].
-Mechanistically, we see that the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward.
-However, as the NL’s policy becomes more deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases.
+-0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs against their opponents and roll out pre-trained meta-policies against newly initialized agents.
+IPD-Rollout: Stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts.
+In this setting, we show that Reciprocator is able to consistently shape an NL over the course of a single learning trajectory with limited rollout samples and observe interesting oscillations in the Reciprocator’s rate of cooperation.
+Figure 2 suggests that these oscillations are driven by the opposing intrinsic and extrinsic rewards, where the derivative of the rate of cooperation corresponds to the reciprocal reward value.
+We provide the following explanation for this phenomenon: although the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward, as the NL’s policy becomes deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases to 0.
 As the extrinsic reward begins to dominate, the Reciprocator essentially reverts into an NL, leading to exploitative defection.
-However, this exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term to become negative, leading to a negative reciprocal reward.
-This, combined with an increase in the NL’s frequency of defection (leading to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$), produces a reversal back to cooperative behavior that is then reinforced by positive reciprocal rewards.
-Figure 3: Shaping an NL in Coins. Percentage of own coins collected by NL during training when facing each opponent (left) and coin counts across 32 steps for Reciprocator vs. NL (right).
+However, exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term, and subsequently the reciprocal reward, to become negative.
+Combined with an increase in the NL’s frequency of defection, which leads to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$, the intrinsic reward produces a reversal back to cooperative behavior that is reinforced by positive reciprocal rewards, and the cycle repeats.
+Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind).
+Total (extrinsic) reward per episode (left), proportion of own coins collected (right).
+Again, Reciprocator and NL-PPO results are plotted on a scale of episodes and MFOS results are plotted on a scale of meta-episodes.
 6.2 Coins
-In this temporally-extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates equal to MFOS.
-This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to this behavior (32,000 steps for Reciprocator vs.
-512,000 steps for MFOS) [Figure 3].
+In this temporally extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates significantly higher than MFOS.
+This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to mutually beneficial behavior (each meta-episode consists of 16 sequential episodes) [Figure 3].
 We show that this change is driven by changes in coin preference rather than in total collection, with both agents collecting at near-optimal pace.
-This is in contrast to MFOS [Appendix B.2], which shapes the NL into collecting fewer coins.
-When two Reciprocators are pitted against each other, we see that they are quickly able to learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play.
-This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieved an average reward of only $\sim 2$ per 32 steps according to Foerster et al.
+This is in contrast to MFOS, which does not shape the NL towards cooperation, but rather uses its pretraining advantage to suppress opponent learning altogether (Khan et al., 2023) by collecting coins faster than the NL can reach them [Appendix B.2].
+When two Reciprocators are pitted against each other, we see that they quickly learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play.
+This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieves an average reward of only $\sim 2$ per 32 steps according to Foerster et al.
 (2018).
-One concern is that the intrinsic reciprocal reward overpowers the extrinsic reward, effectively changing the motivational structure of the game into one that is fully-cooperative.
-To address this, we additionally recorded the cumulative reciprocal reward received in each episode over the course of training [Appendix B.1] to show that the reciprocal reward quickly decreases to a negliglible fraction of the total reward.
-Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind).
-Total reward per episode (left), percentage of own coins collected (right).
-Reciprocators and Naive Learner results are plotted on a scale of single episodes (bottom x-axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode is equivalent to 16 episodes (top x-axis).
+One concern is that the intrinsic reciprocal reward dominates the extrinsic reward, effectively changing the motivational structure of these SSDs into a cooperative game.
+To address this, we recorded the cumulative reciprocal reward received per episode [Appendix B.1] to show that it quickly decreases to a negligible fraction of the total reward, similar to the scale of the reciprocal reward in IPD in Figure 2.
 Together, these results demonstrate that Reciprocators are able to robustly shape the behavior of other agents towards prosocial equilibria during simultaneous learning, achieving state-of-the-art results with fewer assumptions and limitations than existing methods.
@@
 7 Limitations and Future Directions
-Our current implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes.
+Our implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes.
 Using handpicked weights to balance intrinsic and extrinsic rewards opens the door for a suboptimal tradeoff between long-term opponent shaping and short-term return maximization.
-This is partially addressed by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2].
+This is partially mitigated by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2].
 Future work will focus on methods to evaluate reciprocation efficacy across multiple episodes and dynamically tune the balance between reciprocal and extrinsic rewards.
+In terms of evaluation, Khan et al.
+(2023) found that Coins does not require history to enable opponent shaping, since the current state is often indicative of past actions.
+Due to computational limitations, we were unable to assess our method’s performance on the suite of Spatio-Temporal Representations of Matrix Games (STORM) designed to test shaping over longer time horizons.
+We do note that Reciprocators capture both types of memory necessary to achieve shaping as identified by Khan et al.
+(2023): the counterfactual baselines trained on replay buffers serve as a way to capture inter-episode context, and the recurrent policies and influence balance capture intra-episode history.
+Therefore, our method can in theory generalize to STORM environments, although we leave this to future work.
 8 Conclusion and Broader Impacts
@@
 Although these agents exhibit cooperative behaviors when pitted against equivalent opponents, they readily manipulate and exploit agents with simpler learning rules or lower computational capabilities.
-This is a fundamentally undesirable outcome in partially-adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box performance.
-We presented Reciprocators, naïve learning agents equipped with an intrinsic reward for multi-agent settings that seeks to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes.
-We showed that Reciprocators are able to shape the behavior of other agents in the environment over the course of training, and can produce prosocial behaviors in a variety of sequential social dilemmas.
-To the best of our knowledge, Reciprocators comprise the first class of learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or training schedules such as self-play or tracing procedures which control opponent selection and rely on the symmetric player roles.
-We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning algorithms.
-Appendix A Appendix / supplemental material
-All experiments were run on an Nvidia 3070 GPU with 8 GB of VRAM. For IPD-Analytic and Coins experiments involving MFOS agents, we adapted the original code from Lu et al. (2022).
-A.1 Hyperparameter Details
-Each experiment was run 8 times, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins.
+This is a fundamentally undesirable outcome in partially adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box pretrained performance.
+We presented Reciprocators, agents which seek to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes.
+We showed that Reciprocators can both learn prosocial behaviors and induce them from other agents in a variety of sequential social dilemmas, while remaining resistant to exploitation by higher-order agents.
+To the best of our knowledge, Reciprocators represent the first class of reinforcement learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or pretraining routines such as self-play or tracing procedures to control opponent selection.
+We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning opponents.
+Acknowledgements
+This work was supported by the following awards to JCK: National Institutes of Health DP2NS122037 and NSF CAREER 1943467.
+Appendix A Experimental Details
+All experiments were run on Nvidia 3070 GPUs with 8 GB of VRAM.
+Results were averaged over eight random seeds, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins.
+For all experiments involving MFOS agents, we adapted the original code from Lu et al.
+(2022), leaving all architectural choices and hyperparameters the same.
+The architecture for both the actor and the critic consists of a state encoder (two convolutional layers followed by a linear layer, with ReLU activations between layers), followed by hidden linear layers as detailed in the table below and a final output linear layer.
+The actor has a softmax activation to output a policy over the action space, whereas the critic simply outputs a scalar value estimate.
 Hyperparameter IPD-Rollout Coins
-Number of Linear Layers 2 2
+Number of Convolutional Layers - 2
+Convolutional Kernel Size - 3
+Number of Linear Layers 2 1
 Size of Linear Layers 2 16
@@
 PPO-Clip $\epsilon$ 0.1 0.15
-Discount Factor $\lambda$ 0.96 0.99
+Discount Factor $\lambda$ 0.96 0.96
 Entropy Coefficient 0.02 0.01
 Table 2: General PPO parameters.
-With the exception of a linear layer size of 32 for Coins, network parameters to estimate the various target functions to compute the VI are the same as those listed for the PPO network in Table 2.
-Target estimates for IPD policies were computed by averaging across observed choices for each of the five possible states (start, CC, CD, DC, DD).
-Replay buffer sizes are in units of episodes, where each episode consists of a batch of 32 steps.
-Then, for example, a replay buffer size of 4 for Coins corresponds to 32 steps $\times$ 4 episodes $\times$ 2048 batch size = 262,144 steps of experience stored in the replay buffer.
+With the exception of a linear layer size of 32 (instead of 16) for Coins, the network architecture and parameters to estimate the various target functions to compute the $VI$ are identical to those described for the PPO components in Table 2.
+For IPD-Analytic, target estimates for opponent policies in IPD were computed as the frequencies of observed choices for each of the five possible states (start, CC, CD, DC, DD).
+For Coins, we implement the decomposition of $VI$ given in Equation 2, separately predicting the immediate reward and the transition probabilities for each possible next state.
+We collect one batch of episodes at the beginning of each experiment to initialize the influence estimators.
 Hyperparameter IPD-Analytic IPD-Rollout Coins
 Replay Buffer Size (in episodes) 5 1 4
+Target Training Batch Size - - 4096
+Batches per Epoch - - 64
 Target Function Update Period 10 3 1
@@
 Table 3: Reciprocator-specific parameters.
-Again, network parameters for the outer meta-policy and inner actor and critic networks are the same as the general PPO parameters above.
-Hyperparameter IPD-Analytic IPD-Rollout Coins
-Outer Episode Length 16 16 16
-Inner Episode Length - 32 32
-Table 4: MFOS-specific parameters.
+Replay buffer sizes are in units of episodes, where each episode consists of a batch of 32 steps.
+A replay buffer of size 4 for Coins corresponds to 32 steps $\times$ 4 episodes $\times$ 2048 batch size = 262,144 steps of experience.
 Appendix B Additional Experiments
 B.1 Reciprocal Reward in Coins
-An example run of two Reciprocators in Coins showing the relative magnitude of intrinsic reciprocal reward and extrinsic environmental rewards over the course of training.
+A representative run of two Reciprocators in Coins showing the relative magnitude of the intrinsic reciprocal reward and extrinsic environmental rewards over the course of training.
 Figure 5: Mean extrinsic reward (left), mean intrinsic reward (right) for two Reciprocators in Coins.
 B.2 Coin Counts vs. NL-PPO
-We show the total coin counts over the course of training runs vs.
-an NL-PPO in order to show that changes in P(Own Coins) are driven by changes in coin color preference rather than changes in the total number of coins collected.
-Figure 6: Percentage of own coins collected by NL-PPO (left) and total number of coins collected by NL-PPO (right) vs. each baseline in Coins.
+We display the total coin counts of an NL when faced against each other type of agent in order to show that changes in P(Own Coins) when faced by a Reciprocator are driven by changes in coin color preference rather than changes in the total number of coins collected.
+In contrast, we see that the NL collects far fewer coins when faced with an MFOS agent, providing support for Khan et al.
+(2023)’s claim that MFOS suppresses rather than shapes learning.
+Note that this figure is generated from the same experimental data as Figure 3.
+Figure 6: Total number of coins per 32 steps collected by NL-PPO (right) vs. each baseline in Coins.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims made in the abstract and introduction do accurately reflect the paper’s contributions and scope.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We include a Limitations section in the main text.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This paper focuses on empirical evaluation and does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: The paper includes all network parameters, algorithmic details, and hyperparameters necessary to reproduce the main experimental results.
+A zip file containing the code is also provided in the supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Code is linked in the abstract.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Training details are described in the Implementation Details (section 5.3) and hyperparameters are provided in the Appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Number of replicates are provided in the Appendix and error bars (depicting standard error of the mean) are defined in Implementation Details.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Compute resources are described in the Appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Broader impacts of this work are discussed in the Conclusion and Broader Impacts section.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: These models are evaluated on simple game-theoretic environments and do not pose a high risk for misuse.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We use previously released code from the MFOS paper and an adapted open-source implementation of PPO, both of which are credited in the relevant code files and in the Appendix.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: New assets, i.e., code, is well-documented via comments and docstrings and is linked in the abstract.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.